Suppose the weight of full-grown oranges in a citrus grove are normally distributed with a mean of 7.3 ounces and standard deviation of 2.1 ounces. Determine the probablity that 50 oranges randomly picked from the grove have a mean weight greater than 10.5 ounces.
1
Expert's answer
2011-05-19T07:02:43-0400
The sample mean weight X has normal distribution
N(7.3, 2.1/√50)=& N(7.3, 0.297)

Hence
Z=(X-7.3)/0.297
has normal distribution N(0,1).
Therefore
P(X > 10.5) = P(Z > (10.5-7.3)/0.297& ) =& P(Z>10.77)& = 1 - P(Z<10.77) =& 1- F(10.77) =0.99999